OpenStudy (anonymous):

Please help? Precalc? Solve the equation for all real values of x. The equation is (1/(cosx-sinx))=cosx+sinx.

OpenStudy (anonymous):

\[[answer: \pi k]\] but what is the process for getting that?

OpenStudy (anonymous):

\[\cos^2x-\sin^2x=1\]

OpenStudy (raden):

an identity : cos^2 (x) - sin^2 (x) = cos(2x)

OpenStudy (anonymous):

Ohh so then it would be cos(2x)-1=0?

OpenStudy (anonymous):

cos(2x)=1

OpenStudy (raden):

yeah, 1 is cos(0) so, cos(2x) = cos(0) again, use the identity of cos(x) = cos(x + k * 2pi)

OpenStudy (raden):

cos(2x) = cos(0) cos(2x) = cos(0 + k * 2pi) 2x = 0 + k * 2pi 2x = k * 2pi x = k * pi

OpenStudy (anonymous):

Ohh okay, that's what I couldn't figure out. I didn't know cos(0)=1. Thank you!

OpenStudy (raden):

you're welcome